1. Consider the market for yo-yos (sometimes called “Yo Yo Ma”). The supply curve is given by P = 10 + 0.05Q. The demand curve is given by P = 25 – 0.1Q.

  1. Are these inverse, or perverse, supply and demand curves? Why? How do you tell?

  2. Plot them and label the intercepts (on both axes, in the case of the demand curve) and the slopes of each.

  3. Calculate the market equilibrium with no government regulation (P, Q, CS, PS).

2. The Congressional yo-yo committee (“Congressional Y o-Y os”) decides that yo- yo prices move up and down with too much volatility. It therefore drafts a regulation that would impose a price ceiling of $17 per yo-yo.

  1. At the current market equilibrium, does this regulation make sense? Why or why not? Explain.

  2. Are yo-yo makers likely to lobby Congress in favor of this law or against? Explain.

3. After stringing along the yo-yo industry with this proposed regulation for a while, Congress reverses itself and decides to impose a price floor of $17.

  1. Calculate the new industry output that results. Also calculate the shortage (excess demand) or surplus (excess supply) that results. Show your work graphically to illustrate your ideas.

  2. Calculate the new CS and PS that results from the regulation.

  3. Calculate any “transfer” between producers and consumers that is caused by the regulation. Which transfers to the other?

  4. Calculate any DWL caused by the regulation.
Answer #1

Just to clarify: In the problem above, both price ceiling and price floor are at $17, which is above the equilibrium price.

Answer:

Demand curve =  P = 25 – 0.1Q

Supply curve = P = 10 + 0.05Q

The demand curve is a regular curve for normal goods because it is inversely related (negative sign) to the price.

The supply curve is a regular curve with a direct (positive) relation to the price.

To find the equilibrium price and quantity (and therefrom other intercepts), we need to equate both the curves.

P = 25 - 0.1Q = 10 + 0.05Q

That is, 25 - 0.1Q = 10 + 0.05Q

25 - 10 = 0.05Q + 0.1Q

15 = 0.15Q

Therefore, Q = 100

Plugging in value of Q in the equation we can find P.

P = 25 - 0.1Q

P = 25 - 0.1 * 100 = 25 - 10 = 15. P = 15

Verifying for supply, P = 10 + 0.05Q = 10 + 0.05 * 100 = 10 + 5 = 15. P = 15

Plotting the graph for various values of price and quantity:

  Price 30 25 TO Demand 100 150 200 250 300 Quantity

Without market regulation,

P = $ 15, Q = 100 units of yo-yo

Consumer surplus = ½ * 10 * 100 = 500

Producer surplus = ½ * 5 * 100 = 250

2. The regulation to impose a price ceiling does not make sense.

Reason: (i) There is fall in total revenue earned by producers by $140

[ ( 15 * 100) - (17 * 80) = 1500 - 1360 = 140]

(ii) There is fall in consumer surplus (welfare of the people) by 180

[( ½ * 10 * 100) - ( ½ * 8 * 800)

(ii) There is deadweight loss of $ 30.

(½ * 3 * 20 = 30)

Producers will lobby for it because there is increase in producer surplus (incentive for producers to produce more).

The new graph will be:

Price 30 25 Supply Odweg LOSS Demand So 9100 150 200 250 300 Quantity

3. When price floor is set at $17, it becomes the binding price. Producers cannot sell below $17. Therefore, there is excess supply in the market. This is called surplus.

The new graph will be:

Price 30 25 Surplus E Deadweight Log Demand so S 100 150 200 250 300 Quantity

Calculate the new industry output that results. Also calculate the shortage (excess demand) or surplus (excess supply) that results. Show your work graphically to illustrate your ideas.

Ans: New output will be 80 units of yo-yo. (Q1 in the graph).

There will be surplus of 50 units (130 - 80)

Calculate the new CS and PS that results from the regulation.

Ans:

New CS = ½ * 8 * 80 = $ 320

New PS = (old PS - ½ * 1 * 20) + (2 * 80)

= 250 - 10 + 160 = $ 400

Calculate any “transfer” between producers and consumers that is caused by the regulation. Which transfers to the other?

Ans: There has been transfer of welfare from CS to PS. To the extent of (2 * 80) $160

Calculate any DWL caused by the regulation.

Ans: Deadweight loss will be the shaded area in the graph. ½ * 3 * 20 = $ 30